narutouzamaki2:

Gorillas have 48 chromosomes. The egg cell of a female gorilla has chromosomes. 24 96 48

Hoodmemes:

\[48\div 2= (blank)\]

ramen:

The reason why is because half of the chromosomes come from one parent and the other half comes from the other parent.

ramen:

so yea 48/2 = 2

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen so yea 48/2 = 2 \(\color{#0cbb34}{\text{End of Quote}}\) re check that

ramen:

huh?

Hoodmemes:

u said 48/2=2 incorrect statment

ramen:

oh wait oops i didn't mean to put the 2 on the end lol

ramen:

That was an accident i meant to leave the 2 out

Hoodmemes:

oh

ramen:

im not stupidoto

Hoodmemes:

I take things very literal srry.

ramen:

It's alright lmao I do too

narutouzamaki2:

so its 24

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @narutouzamaki2 so its 24 \(\color{#0cbb34}{\text{End of Quote}}\) correct-o-predna

ramen:

yes

narutouzamaki2:

thx

Hoodmemes:

np

ramen:

fasho